Math. 312, Sec. 500 Fall, 2000

Topics in Applied Mathematics II (partial differential equations)

Classes: MWF 3:00–3:50, Blocker 164

Instructor: S. A. Fulling Blocker 620H 845-2237 fulling@math.tamu.edu If I am not in my office, you can leave a note in my mailbox (in the room op- posite the math department office, 6th floor of Blocker) or in the plastic pouch beside my office door.

Temporary office hours: MF 11:00–12:00, R 3:00–3:50. Permanent office hours will be announced later.

Prerequisite: M. 308 or 451 (differential equations). (M. 311 or other linear algebra will help, but is not required.)

Textbooks:

1. required: R. Haberman, Elementary Applied Partial Differential Equations with Fourier Series and Boundary Value Problems, 3rd edition, Prentice–Hall, 1999.

2. strongly recommended: S. A. Fulling, Math 312 lecture notes, for sale at Copy Corner (1404 Texas Ave. S.), Packet #162. There is a separate booklet of old tests, Packet #163, with a disclaimer that they do not precisely match the present syllabus.

3. optional, cheap: Evans Harrell and James Herod, Linear Methods of Applied Mathematics, http://www.mathphysics.com/pde/ . (Read the copyright notice!)

4. optional, cheap: M. R. Spiegel (Schaum’s Outline Series), Fourier Analysis.

5. optional, cheap: G. P. Tolstov, Fourier Series, Dover.

Grading system: Hour tests:  100 × 3 = 300 Final exam:  200 Homework and class participation:  \underline{200} Total  700

The “curve” will be at least as generous as the “standard” scale [i.e., 90% (= 630 pts) will guarantee an A, etc.].

Dates of hour tests: Sept. 22, Oct. 20, Nov. 17 (Fridays) Final exam: Tuesday, Dec. 12, 10:30–12:30